OpenStudy (anonymous):

2(3x+5)=46 A)4 B)5 C)6 D)7

OpenStudy (lυἶცἶ0210):

Try distributing: \(\large (2)(3x)+(2)(5)=46\)

syed98 (syedmohammed98):

6x + 10 = 46 6x = 46 - 10 6x = 36 x=36/6 x = answer =)

OpenStudy (zale101):